14. A fly, buzzing about the room, calculates the angle of depression of the base of an 8-foot wall is 33.7° and the angle of elevation of the top of the wall is 12.5°. Find to the nearest foot the horizontal distance the fly is from the wall.
1
Expert's answer
2020-03-06T13:37:10-0500


Here as per question the height of wall is 8ft

so , i draw the above diagram as per question requirement

Let x be the distance of CD, then DO will be 8-x

Now, In "\Delta CDB, CD =x , BD" is the base ,

so, "\tan \theta = \frac{CD}{BD},"


"\tan (12.5^o) = \frac{x}{BD}"


BD"= \frac{x}{.2217}" ........................1

similarly , In "\Delta BDO"


"\tan (33.7^o)=\frac{8-x}{BD}"


BD= "\frac{8-x}{.6669}" .......................2


From equation 1 and 2 we can write


"\frac{x}{.2217} =\frac{8-x}{.6669}"


".6669x= 1.7736-.2217 x"

."0.8907x=1.7736"

"x= 1.99 ft"

so just we put this value in equation 1 we get

BD = 8.98ft

so ,the dustance of fly from wall is 8.98 ft we can take it approx as 9ft